OpenStudy (angel_kitty12):

Also this problem please help ASAP 8^-1/2

OpenStudy (anonymous):

0.35

OpenStudy (alexandervonhumboldt2):

\[8^{-1/2}=\frac{ 1 }{ 8^{1/2} }=\frac{ 1 }{ \sqrt{8} }=\frac{ 1 }{ 2\sqrt{2} }\]

OpenStudy (angel_kitty12):

But how?

OpenStudy (alexandervonhumboldt2):

i showed you @Angel_Kitty12

OpenStudy (angel_kitty12):

Oh. But isn't it not allowed to have a radical in the denominator?

OpenStudy (alexandervonhumboldt2):

easy: \[\frac{ 1 }{ \sqrt{8} }=\frac{ 1 }{ \sqrt{8} }*\frac{ \sqrt{8} }{ \sqrt{8} }=\frac{ \sqrt{8} }{ 8}\]

OpenStudy (angel_kitty12):

So would it be 2sqrt2/8 ?

OpenStudy (alexandervonhumboldt2):

\[\frac{ \sqrt8 }{ 8}=\frac{ 2\sqrt{2} }{ 8 }\]

OpenStudy (alexandervonhumboldt2):

so yeah you are correct

OpenStudy (angel_kitty12):

Then simplifies to sqrt2/4 ?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (alexandervonhumboldt2):

\[\frac{ \sqrt2 }{ 4 }\]

OpenStudy (angel_kitty12):

And that's the answer right?

OpenStudy (alexandervonhumboldt2):

yeah